# DMol: A Schedule-Driven Diffusion Model for Highly Efficient and Versatile Molecule Generation

### v1

@@ E.3 Selection of the parameter $k$
[…]
@@
Table 7: DMol with different values of $k$ on MOSES dataset
E.4 Comparison with GruM
We also compare CDMol to GruM (Jo et al., 2024) on the 2D molecule generation task. Since GruM was designed for small-scale datasets, we ran both methods on QM9 to ensure fairness. As shown in Table 8, GruM achieves slightly higher validity but significantly lower uniqueness and novelty scores. Moreover, GruM does not scale well, and we were unable to run it to completion on Guacamol using hardware that successfully supports both DiGress and DMol.
[…]
GruM 99.4 85.6 26.1 85.2 22.2
CDMol 98.3 97.7 78.2 96.0 75.1
Table 8: Comparison between CDMol and GruM on QM9 dataset
Appendix F Examples of Generated Molecules.
Figure 8 shows the sampling process.
Figure 9, 10, 11 show examples of molecule generation.
@@ E.4 Comparison with GruM
Appendix G Docking Results
We used AutoDock Vina version 1.2.0 (Goodsell & Olson, 1990) to visualize the docking of two of the top-scoring generated molecules (with respect to the ChEMBL likeness) that also exhibit strong biding affinities on the 1PXH and 4MQS proteins when compared to reference drug ligants. The 4MQS complex represents the active human M2 muscarinic acetylcholine receptor bound to the antagonist iperoxo, while 1PXH represents the protein tyrosine phosphatase 1B with bidentate inhibitor compound 2. The free energy results are shown in Figures 12 and 13, and they indicate that the generated molecules can signficantly outperform or underperform existing reference drugs.
(a) Reference: Free energy = $-8.7$ kcal/mol
(b) Generated drug 1: Free energy = $-6.4$ kcal/mol
(c) Generated drug 2: Free energy = $-6.0$ kcal/mol
@@ Appendix G Docking Results
(a) Reference: Free energy $-6.4$ kcal/mol
(b) Generated drug 1: Free energy = $-6.6$ kcal/mol
(c) Generated drug 2: Free energy = $-9.1$ kcal/mol
Figure 13: The reference molecule (a) and two of our generated molecules (b) and (c) docked at the protein receptor 4MQS, and the corresponding free energies. The generated molecules significantly outperform with respect to the reference, and the sizes/masses of the drug molecules are comparable. This further confirms the importance of focusing on the right molecular weight and size of the generated molecules.

### v3

@@ G.3 Selection of the parameter $k$
Table 7: DMol with different values of $k$ on MOSES dataset
$k$ Validity ↑ Uniqueness ↑ Novelty ↑ Sample Time ↓
@@
G.4 Complete Experimental Results for the QM9 Dataset
For complete performance comparison on the QM9 dataset, please refer to Table 8.
Table 8: Performance comparison on the QM9 dataset.
Model V↑ V.U.↑ V.U.N.↑
GraphVAE 56.1 42.8 26.5
GT-VAE 74.8 16.5 15.6
Set2GraphVAE 60.0 56.8 -
SPECTRE 87.5 31.4 29.0
GraphNVP 83.5 18.4 -
GDSS 96.0 94.6 -
GruM 99.4 85.1 22.2
DiGress 97.8 95.2 31.8
DisCo 98.2 95.6 57.5
DeFoG 97.9 95.9 62.8
DMol(Ours) 98.3 96.0 75.1
G.5 Forward Process Schedule
To examine the impact of the cosine schedule on different molecular scaffolds and topologies, we conducted an ablation study comparing different values of the hyperparameter $c$ in the cosine schedule $\alpha=\cos^{2}(0.5\pi(t/T+c)/(1+c))$, as well as comparing cosine versus linear schedules on the MOSES dataset. The results are presented in Table 9.
Table 9: Ablation study for forward process schedules on the MOSES dataset. The cosine schedule with $c=0.008$ achieves optimal performance across all metrics.
Method Validity↑ Uniqueness↑ Novelty↑ Filters↑ FCD↓ SNN↑ Scaf↑ ChEMBL↑ QED↑
cos ($c=0.004$) 86.9 100 100 97.4 1.13 0.57 14.2 4.5022 0.8022
cos ($c=0.006$) 87.5 100 100 97.7 1.12 0.57 14.6 4.5020 0.8037
cos ($c=0.008$) 87.8 100 100 97.8 1.12 0.58 14.8 4.5033 0.8055
cos ($c=0.01$) 87.3 100 100 97.0 1.19 0.55 14.3 4.5018 0.8042
Linear 85.2 99.8 99.9 96.1 1.19 0.56 14.2 4.5011 0.8016
The experimental results demonstrate that varying the value of $c$ has a relatively small impact on crucial performance metrics, with an optimal performance achieved at $c=0.008$, which is the setting used in our experiments. This finding aligns with the findings by DiGress, which also adopts $c=0.008$ as the default value.
When comparing cosine and linear schedules, we see that the cosine schedule clearly outperforms the linear schedule. This advantage stems from the fact that the cosine schedule adds less noise per timestep during both the initial and final phases of the diffusion process, compared to intermediate stages. This distribution better aligns with the learning characteristics of diffusion models. Specifically, maintaining relatively small noise levels during the early and final diffusion phases facilitates the model’s learning of fine-grained data features and ensures “smoother” convergence. In contrast, the uniform noise distribution in linear schedules may prove either overly aggressive or overly conservative at certain stages, leading to compromised training efficiency and performance.
G.6 Scaffold-Constrained and Conditional Generation
To avoid issues associated with direct conditional generation, we instead proposed a scaffold-informed pipeline. The key ideas are to cluster the training set molecules according to either their RDKit features, or embeddings generated by Mol2Vec or Grover Jaeger et al. [2018], Rong et al. [2020], and then run DMol on the samples in sufficiently large cluster groups separately. More details are provided below.
RDKit-Based Scaffold Classification: Molecules are categorized into four scaffold classes following RDKit recommendations: aromatic monocyclic, aromatic monocyclic heterocycle, fused bicyclic, and aromatic heterocyclic.
GNN, Transformer and Other Embedding Classification Methods: One can use Mol2Vec or Grover to perform molecular graph embeddings, and follow up with Kmeans++ clustering.
The detailed analysis of the pros and cons of this approach compared to classical conditional generation will be discussed in more detail in a companion paper.
G.7 Motif Probability Distribution Analysis
Table 10 presents the complete per-motif probabilities for the top 30 most frequent motifs in the MOSES dataset. Table 11 shows the marginal distributions of node and supernode categories.
Table 10: Comparison of motif probabilities between training and generated sets on the MOSES dataset. IDs 0140-14 correspond to supernodes, while IDs 152915-29 are nonsupernode motifs.
ID Motif (SMILES) Prob on Training Set Prob on Generated Set
Supernode Motifs (IDs 0140-14)
0 c1ccccc1 0.7753 0.7861
1 c1ccncc1 0.1829 0.1630
2 c1cnnc1 0.1019 0.0912
3 C1CCNCC1 0.0852 0.0783
4 C1CCNC1 0.0798 0.0662
5 c1cscc1 0.0763 0.0740
6 c1ccsn1 0.0715 0.0686
7 C1COCCN1 0.0625 0.0651
8 C1CNCCN1 0.0620 0.0681
9 c1ccoc1 0.0573 0.0429
10 c1cncnc1 0.0559 0.0512
11 c1cncn1 0.0464 0.0416
12 c1ncon1 0.0459 0.0458
13 c1ncnn1 0.0453 0.0415
14 C1CCCCC1 0.0351 0.0301
Nonsupernode Motifs (IDs 15-29)
15 C1CC1 0.0298 0.0242
16 c1cnoc1 0.0289 0.0238
17 C1CCCC1 0.0246 0.0281
18 C1CCOC1 0.0231 0.0241
19 c1ccnc1 0.0214 0.0191
20 c1cCNCC1 0.0202 0.0169
21 c1ccnnc1 0.0191 0.0123
22 c1nnnn1 0.0177 0.0161
23 c1nncs1 0.0164 0.0143
24 c1cnccn1 0.0154 0.0127
25 c1cnnn1 0.0151 0.0104
26 c1cocn1 0.0146 0.0088
27 c1nnco1 0.0135 0.0090
28 c1cCCCC1 0.0123 0.0098
29 c1cNCC1 0.0114 0.0102
Average Probabilities
Top 15 (Supernodes) mean 0.1189 0.1143
Last 15 (Nonsupernodes) mean 0.0189 0.0160
Table 11: Marginal distributions of node and supernode categories in the MOSES dataset. Supernodes are treated as new node types alongside individual atoms.
Node / Supernode Type Training Set Generated Set
Individual Atoms
C 0.4897 0.4940
N 0.1511 0.1506
S 0.0170 0.0167
O 0.1654 0.1644
F 0.0246 0.0243
Cl 0.0093 0.0091
Br 0.0025 0.0026
Supernode Motifs
c1ccccc1 0.0736 0.0750
c1ccncc1 0.0132 0.0130
c1cnnc1 0.0081 0.0077
C1CCNCC1 0.0068 0.0070
C1CCNC1 0.0049 0.0045
c1cscc1 0.0055 0.0052
c1ccsn1 0.0056 0.0050
C1COCCN1 0.0035 0.0036
C1CNCCN1 0.0036 0.0031
c1ccoc1 0.0041 0.0039
c1cncnc1 0.0044 0.0040
c1cncn1 0.0010 0.0008
c1ncon1 0.0027 0.0025
c1ncnn1 0.0024 0.0020
C1CCCCC1 0.0010 0.0008
Appendix H Examples of Generated Molecules.
Figure 8 illustrates the sampling process.
Figure 9, 10, 11 shows examples of generated molecules.
@@ G.7 Motif Probability Distribution Analysis
Appendix I Sample Docking Results
We used AutoDock Vina version 1.2.0 Goodsell and Olson [1990] to visualize the docking of two of the top-scoring generated molecules (with respect to the ChEMBL likeness) that also exhibit strong biding affinities on the 1PXH and 4MQS proteins when compared to reference drug ligants (the reason behind the selection of these proteins is confidential information). The 4MQS complex represents the active human M2 muscarinic acetylcholine receptor bound to the antagonist iperoxo, while 1PXH represents the protein tyrosine phosphatase 1B with bidentate inhibitor compound 2. The free energy results are shown in Figures 12 and 13, and they indicate that the generated molecules can significantly outperform or underperform existing reference drugs.
(a) Reference: Free energy = 8.7-8.7 kcal/mol
(b) Generated drug 1: Free energy = 6.4-6.4 kcal/mol
(c) Generated drug 2: Free energy = 6.0-6.0 kcal/mol
@@ Appendix I Sample Docking Results
(a) Reference: Free energy = 6.4-6.4 kcal/mol
(b) Generated drug 1: Free energy = 6.6-6.6 kcal/mol
(c) Generated drug 2: Free energy = 9.1-9.1 kcal/mol
Figure 13: The reference molecule (a) and two of our generated molecules (b) and (c) docked at the protein receptor 4MQS, and the corresponding free energies. The generated molecules significantly outperform with respect to the reference, and the sizes/masses of the drug molecules are comparable. This confirms the importance of focusing on the right molecular weight/size.
Appendix J Broader Impact
This work presents DMol, an efficient molecular graph generation framework with a strong potential for practical applications in accelerating drug discovery, designing novel materials, and advancing computational chemistry. By substantially reducing computational requirements while maintaining or improving molecular quality metrics, our approach could democratize access to molecular design tools, enabling researchers with limited computational resources to engage in this field. The efficiency gains may also enable exploration of larger chemical spaces, potentially leading to discoveries that address pressing challenges in healthcare, energy storage, and environmental remediation. However, we acknowledge that molecular generation technologies could potentially be misused for designing harmful substances if deployed without appropriate safeguards. Additionally, as with many other AI systems, there is risk of amplifying biases present in training data, which could limit the diversity of generated molecules or reproduce historical biases in pharmaceutical development. We encourage future work to address these concerns through development of responsible use protocols, integration of toxicity and safety prediction tools, and careful curation of training datasets to ensure equitable representation across chemical domains. The machine learning community should collaborate with domain experts in chemistry, biology, and ethics to ensure that advances in molecular generation technologies maximize societal benefit while minimizing potential harms.
Appendix K Limitations
Despite the many documented advantages of DMol in terms of efficiency and molecular quality, several limitations remain. First, our approach still struggles with generating certain complex motif structures and stereochemistries, which are crucial for therapeutic applications. Second, while we achieve high validity scores, we do not explicitly enforce chemical constraints during generation, occasionally producing molecules that are formally valid but synthetically impossible to bring into existence. Third, our evaluation focuses primarily on small drug-like molecules; the performance on larger biomolecules, polymers, or metal-organic frameworks remains unexplored. Finally, the model’s adaptability to conditional generation tasks (e.g., optimizing for specific target properties) requires additional architectural modifications that may affect the established efficiency gains. Future work should address these limitations to further bridge the gap between computational generation and practical chemical synthesis. Most importantly, this and all other related works should put more effort in identifying more comprehensive evaluation metrics for the generated molecules, since an overwhelming number of created samples cannot be synthesized or do not dock on any known protein.
Appendix L Future Work
Future work on DMol will focus on extending the model to conditional molecule generation, where specific chemical properties can be targeted through controlled diffusion processes. We also plan to address the remaining stereochemistry challenges by incorporating 3D structural information into the diffusion process. Additionally, we aim to develop more sophisticated motif compression strategies that can handle a wider variety of molecular scaffolds while preserving chemical feasibility. Exploring the application of DMol to larger biomolecular structures and integrating it with downstream property prediction models presents promising research directions. Finally, we intend to investigate dynamic scheduling of diffusion steps based on molecular complexity to further optimize computational efficiency.